OpenStudy (anonymous):

How do I simplify 2m/m-9 + 18/9-m?

OpenStudy (anonymous):

2m 18 ----- + ------ m - 9 -(m-9)

OpenStudy (eyust707):

first muliply the right side by -1/-1

OpenStudy (eyust707):

then add

OpenStudy (anonymous):

really?

OpenStudy (eyust707):

factor out a 2

OpenStudy (anonymous):

'cause my teacher said: 9-m = --9 +- m = - (-9 + m) = - (m-9)

OpenStudy (eyust707):

yeah cause we want a common denominator

OpenStudy (eyust707):

ohh i see

OpenStudy (eyust707):

the second post was your attempt?

OpenStudy (eyust707):

yea so you got it

OpenStudy (anonymous):

no I don't// cos I don't know what to do after..

OpenStudy (eyust707):

2m 18 (-1) ----- + ------ * ---- m - 9 -(m-9) (-1)

OpenStudy (eyust707):

that bring the negitive up top

OpenStudy (anonymous):

I only got to 2m 18 ---- + ---- m-9 -(m-9)

OpenStudy (anonymous):

OHHH.

OpenStudy (anonymous):

why did you multiply by -1?

OpenStudy (eyust707):

which is also the same as 2m 18 ----- - ------ m - 9 (m-9)

OpenStudy (eyust707):

cause you can mupliply any thing by -1/-1

OpenStudy (eyust707):

without changing it

OpenStudy (eyust707):

a/a = 1

OpenStudy (eyust707):

i basically just muplitplied it by -1/-1 = 1

OpenStudy (eyust707):

from there you have a common denominator

OpenStudy (eyust707):

and can simplify

OpenStudy (anonymous):

ohhhh, i see. Thank you!

OpenStudy (eyust707):

no problem =P

OpenStudy (phi):

the final answer is 2

OpenStudy (anonymous):

^ why is it 2?

OpenStudy (phi):

how far did you get?

OpenStudy (anonymous):

because I have 2m-18 ------ m-9

OpenStudy (phi):

you can factor 2 out of each term in the numerator to get 2(m-9)

OpenStudy (phi):

and (m-9)/(m-9) =1

OpenStudy (anonymous):

what do you mean factor 2 out of each term?

OpenStudy (phi):

you know 2x+2y= 2(x+y), for example?

OpenStudy (phi):

or 2m-18= 2*m- 2*9 = 2(m-9)

OpenStudy (anonymous):

ohhh, yup.

OpenStudy (anonymous):

and what do you mean by (m-9)/(m-9)? Is it possible for you to draw how you did it?

OpenStudy (phi):

you know a number divided by itself is 1 (m-9) is a number (you just don't know exactly what number)

OpenStudy (anonymous):

ohh, yup.